### v4

# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

Abstract
This research addresses the challenge of integrating forecasting and optimization in energy management systems, focusing on the impacts of switching costs, forecast accuracy, and stability. It proposes a novel framework for analyzing online optimization problems with switching costs and enabled by deterministic and probabilistic forecasts. Through empirical evaluation and theoretical analysis, the research reveals the balance between forecast accuracy, stability, and switching costs in shaping policy performance. Conducted in the context of battery scheduling within energy management applications, it introduces a metric for evaluating probabilistic forecast stability and examines the effects of forecast accuracy and stability on optimization outcomes using the real-world case of the Citylearn 2022 competition. Findings indicate that switching costs significantly influence the trade-off between forecast accuracy and stability, highlighting the importance of integrated systems that enable collaboration between forecasting and operational units for improved decision-making. The study shows that committing to a policy for longer periods can be advantageous over frequent updates. Results also show a correlation between forecast stability and policy performance, suggesting that stable forecasts can mitigate switching costs. The proposed framework provides valuable insights for energy sector decision-makers and forecast practitioners when designing the operation of an energy management system.
keywords:
Energy Management Systems, Forecasting, Optimization, Switching Costs, Forecast Stability, Model Predictive Control
{highlights}
The paper evaluates the impact of forecast stability and accuracy on the downstream performance of energy management systems under switching costs.
A novel metric, Scenario Distribution Change (SDC), is proposed to assess the stability of probabilistic scenario sets, extending the Wasserstein distance.
Forecast accuracy and stability are shown to influence the optimal policy commitment level. There is a theoretical trade-off between switching costs and forecast errors.
Empirical analysis is conducted using the CityLearn 2022 competition case study, highlighting the trade-offs between forecast updates, stability, and operational outcomes. The longer than minimal commitment periods are found to improve system performance in case of stochastic optimization with probabilistic scenario sets.
1 Introduction
Managing energy assets within a grid system presents a challenging task characterized by decision-making under uncertainty. The inherent dynamics and stochasticity of the environment make this a complex system, constantly evolving and demanding decisions to be made with incomplete knowledge of the future over limited time windows, a scenario typically addressed through online optimization.
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9, 10].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information. Striking a balance between forecast accuracy and the impact of switching costs is crucial. This paper explores how this balance shapes system design and informs operational strategies. Consequently, this research aims to answer the following questions:
  • 1.
    Do switching costs influence the balance between forecast accuracy and stability?
  • 2.
    What are the optimal strategies for forecast revision in the presence of such costs?
  • 3.
    How can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
1.1 Related work
Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’. Such approach is proposed in Elmachtoub and Grigas [11], with optimization cost integrated directly into the forecast loss function to align forecasts with end-use optimization. This optimizes the prediction model in the forecasting stage for better decision-making in the optimization stage. The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges. There are also other works that design an alternative task-specific loss function, such as reported in Mandi et al. [12], and tested on a real-world energy management problem.
Indirect methods regard the two subproblems as separate, also referred in literature as ‘predict, then optimize’. This approach is regarded as standard and is widely applied in practice. In the study by Vanderschueren et al. [13], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification. Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of the classifier outperforms the integrated approach. Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [14], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task. Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment. A relevant study by Prat et al. [15] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach. The study introduces a verifiable condition to check if a chosen planning horizon is sufficiently long. Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The frequency of retraining is studied by Spiliotis and Petropoulos [16], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
We are particularly interested in studying the model deployment frequency. More specifically, we look at the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
(a) FHC
(b) RHC
Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control
The research conducted by Chen et al. [17] highlights the challenges posed by noisy predictions in designing Online Convex Optimization (OCO) algorithms. Several solutions, such as Averaging Fixed Horizon Control (AFHC) and its generalized counterpart, Committed Horizon Control (CHC) [18], propose aggregation strategies over multiple revisions of plans to achieve sublinear regret and improve robustness to prediction errors. Similarly, the Feasible Fixed Horizon Control (FFHC) algorithm [19] incorporates a regularization term to handle multi-interval optimization problems while ensuring feasibility and robustness against ramping costs.
While these approaches demonstrate the benefits of aggregating policies over prediction updates, the current work shifts focus to understanding the effects of commitment on forecast stability and downstream optimization in the presence of switching costs. Unlike these aggregation-based methods, this study seeks to explore how the commitment level impacts system performance, with a particular emphasis on the relationship between forecast stability and switching costs.
Forecast stability has been recognized as a critical factor in planning. Building on the comprehensive framework by Godahewa et al. [20], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems.
Unlike supply chain management, where freezing intervals [21, 22] and smoothing techniques [23] have been widely used to address system nervousness and the bullwhip effect [24], energy systems require more dynamic and frequent forecast updates. Switching costs are also observed in energy management systems, such as ramping costs for generating units [7, 25] or toggling server load balancing [8, 26]. However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs. By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
1.2 Contents and Contributions
The principle contributions are as follows:
  • 1.
    First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs. We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
  • 2.
    The effect of forecast accuracy and stability on the downstream performance of the optimization policy is analyzed. We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
  • 3.
    We conduct an empirical evaluation of the stability of point forecasts and probabilistic scenario sets, using deterministic and stochastic optimization approaches, respectively. We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy. We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
  • 4.
    Lastly, the paper concludes with a discussion of the general implications of the results for the design of the energy management systems followed by suggested future research directions.
2 Methodology
2.1 Problem statement
Figure 2: Grid diagram illustrating the coordinated battery scheduling problem.
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [27] and based on a real-world dataset from a grid-interactive community. The CityLearn Challenge is an annual competition that focuses on the development of reinforcement learning agents for energy management in buildings. Participants are provided with a simulated environment that mimics real-world energy consumption and generation patterns. The goal is to optimize the operation of energy systems, such as batteries and HVAC systems, to minimize costs and emissions while maintaining comfort levels.
The series of CityLearn competitions, starting from 2020, was originally set up only for Reinforcement Learning (RL) solutions, where an agent learns to make decisions by performing actions and receiving feedback from the environment. However, from 2022 onwards, alternative methods such as Model Predictive Control (MPC), which uses a model to predict the future behavior of a system and optimize the control actions accordingly, and Rule-Based Control (RBC), which involves using predefined rules to make decisions based on the current state of the system, have also been employed. The competition provides a platform and the open-source simulation environment [^1] https://github.com/intelligent-environments-lab/CityLearn for implementing the operation of energy management systems with setups that include various energy assets, such as solar panels, batteries, heat pumps, connections to the grid, and electric vehicles.
The setup of the investigated problem is illustrated in Figure Figure 2. In general terms, we consider an online optimization problem with switching costs. The data of the competition has the hourly resolution. The core problem is multi-agent scheduling of energy storage over a period of a year. While the problem can be formulated similarly to RL, or a dynamic programming approach, in this paper, we particularly investigate the mechanics with a rolling-horizon approach. This means that the problem is solved iteratively over a planning horizon $H$. However, it is typical for only first $v$ steps to be implemented with the rest of the prediction horizon $H-v$ to be advisory, since the optimization is revised with a period of $v$ steps. The full details of the objective score and relevant constraints are provided in A.1. The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [28], we use the following formulation:
$X_{t}^{FHC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
$\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
$+\beta\left\lVert\tilde{y}_{tt^{\prime}}-\tilde{y}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
In this formulation:
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$. This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
    $\mathbb{C}(\textbf{x};\textbf{S})$ is the cost function, which is a function of actions $\textbf{x}=x_{t},...x_{t+H}$ and states $\textbf{S}=S_{t},...S_{t+H}$.
  • 3.
    $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process. We assume that the weight factors are equal for all scenarios. In the deterministic case, the number of scenarios is equal to 1.
  • 4.
    $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ Represents the cost of the set of decisions $x_{t+1},...,x_{t+H}$ given the future forecast scenarios generated by the model $\theta^{LA}_{i,tt^{\prime}}$
  • 5.
[…]
Optimized battery schedules are evaluated using a set of Key Performance Indicators (KPIs). The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. When grid-related metrics are included, the average includes the grid score $D$, which is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, the combined optimization scores are defined with and without the grid score:
Average Score $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
Average Score (with switching costs) $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$ (3)
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in A.2. Furthermore, In A.3, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. Following the line of proof in Chen et al. [18], we derive the upper bound on the competitive difference between the FHC algorithm and the optimal policy. The analysis is conducted for the deterministic case, where the forecast is a point forecast, then extended to the probabilistic case, where the forecast is represented as a set of scenarios. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
2.2 Optimization
The problem is solved using the Fixed Horizon Control (FHC) algorithm, which is a configuration of model predictive control (MPC) with a fixed commitment horizon. The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps. An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$. Every $v$ time steps, the optimization is rerun to generate a new plan. This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
In the context of the online optimization algorithms, the term “commitment” refers to the duration or number of time steps for which the optimized policy is committed to before the optimization process is rerun to generate a new plan. This concept plays a crucial role in determining how frequently the optimization is updated based on new information, accumulation of forecast errors and changes in the system’s state. An FHC algorithm with the shortest commitment is the Receding Horizon Control (RHC) algorithm. In RHC, the commitment is $v=1$, indicating that the optimization is rerun at every time step.
Similar to the optimization, the forecast is also run in a rolling origin fashion. The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps. In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps. Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. The resulting decision matrix is shown in Table 1.
Table 1: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
Optimization Decision
Forecasting Decision Reuse Revise
Reuse Continue with current plan Evaluate
& adjust current plan
Revise Evaluate &
retain current plan
Re-evaluate both
forecasting & plan
2.3 Forecast
The target variable for forecasting is the aggregate uncontrollable load of the energy community in Figure 2. Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model. The gradient boosting implementation is LightGBM [29]. Predictions are made directly for the next 24 hours with 24 independent models. The scenario sets are generated with added gaussian noise. The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al. [27] and the code repository of the winner solution is available on GitLab [30]. The evaluation conducted in this study is available in the code repository [31].
While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings. This setup follows the problem setting given in the Citylearn Challenge 2022 [27]. Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
2.3.1 Rolling Origin Window
In practical applications, the forecast is issued periodically for a finite number of future time steps. This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps. A schematic of this process is shown in Figure 3. This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration. For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
Figure 3: Updating Rolling Origin Window. The forecast
is updated every $v_{f}$ time steps and possible overlaps between prediction windows are shown.
The rolling horizon setup enables an update of the prediction using the most recent information. When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions. The resulting overlapping interval is highlighted in light grey in Figure 3. These overlaps are considered when evaluating stability of the forecast.
2.3.2 Stability
Stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed in [20], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 4.
(a) Vertical Stability
(b) Horizontal Stability
Figure 4: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions. The vertical stability is the difference between the predictions produced at different origins for the same time period. If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal. Horizontal stability characterizes the difference between predictions generated at the same origin. The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
Similarly to the evaluation in [32], a stability metric is measured for point forecasts using the mean absolute change (MAC). The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the change of predictions within the horizon window generated at the same origin — ${MAC}_{H}$
$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (4)
$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (5)
where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
2.3.3 Stability of Probabilistic Forecasts
In stochastic optimization, scenario stability measures the consistency of probabilistic forecast updates over time. We propose the Scenario Distribution Change (SDC) metric, adapting the Wasserstein Distance, also referred to as Earth Mover Distance, to assess the stability of scenario sets. SDC quantifies the average change between successive scenario updates, providing a clear measure of stability for probabilistic forecasts in dynamic systems.
The Wasserstein distance formulation is symmetric, adheres to the triangle inequality, and effectively compares probabilistic scenario sets. Unlike metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon distance, SDC holds for non-overlapping distributions, making it particularly relevant for applications requiring forecast stability. By incorporating the dimension of the metric space, SDC is conceptually analogous to the Mean Absolute Change (MAC) used for point forecasts.
To assess stability across time, SDC evaluates both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates). For $N$ number of scenarios, these are defined as:
Horizontal Stability (SDCH):
$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (6)
Vertical Stability (SDCV):
$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (7)
2.3.4 Accuracy of Determinstic Forecasts
For point forecasts, the MAE is utilized to quantify the forecast accuracy. It calculates the average absolute difference between each forecasted value and the corresponding actual value. Mathematically, it is expressed as:
$\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (8)
where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
2.3.5 Accuracy of Probabilistic Forecasts
For probabilistic forecasts, we employ the Energy Score (ES). The ES is a multivariate generalization of the continuous ranked probability score (CRPS), a widely used metric for evaluating probabilistic forecasts. The ES measures the distance between the forecasted distribution and the actual value. The lower the ES, the better the forecast. The ES is defined as:
$ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (9)
where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm. The second term in the equation is the mean pairwise distance between the scenarios.
3 Experimentation and Results
Compared to theoretical analysis, practical setups are more complex. They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization. This complexity is due to the necessity for cooperative behavior across multiple buildings. Additionally, various temporal dependencies emerge from switching costs and system feedback. To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
According to the decision matrix in Figure 1, at every time step the forecast and optimization plans can be updated or retained. Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$, respectively. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment. The result is shown in Tables 2(a) and 2(b). The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
(a) deterministic FHC algorithm with point forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.899
2 0.910 0.904
3 0.911 0.907 0.904
4 0.912 0.908 0.907 0.905
5 0.913 0.910 0.908 0.908 0.903
6 0.912 0.910 0.907 0.908 0.908 0.906
7 0.912 0.910 0.907 0.910 0.908 0.909 0.902
8 0.910 0.907 0.908 0.905 0.906 0.908 0.902 0.903
9 0.911 0.909 0.906 0.908 0.908 0.908 0.904 0.908 0.903
10 0.913 0.909 0.909 0.907 0.906 0.908 0.903 0.906 0.905 0.905
11 0.912 0.910 0.909 0.910 0.908 0.909 0.905 0.906 0.906 0.907 0.903
12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
(b) stochastic FHC algorithm with probabilistic forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.875
2 0.888 0.875
3 0.894 0.886 0.873
4 0.895 0.885 0.882 0.874
5 0.899 0.892 0.885 0.886 0.875
6 0.901 0.891 0.881 0.883 0.884 0.876
7 0.901 0.896 0.887 0.888 0.886 0.883 0.875
8 0.900 0.894 0.887 0.884 0.884 0.885 0.882 0.874
9 0.902 0.894 0.883 0.887 0.884 0.882 0.883 0.880 0.876
10 0.903 0.894 0.888 0.887 0.881 0.884 0.883 0.880 0.883 0.876
11 0.902 0.895 0.889 0.887 0.885 0.885 0.884 0.880 0.884 0.882 0.876
12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
Table 2: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.4. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours. For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [27]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period. This observation is notable as the most expected behavior is that best performance is achieved with shortest revision periods for both forecast types, as more up-to-date forecasts are provided.
These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
Accuracy
Stability: vertical (top) and horizontal (bottom)
Optimization Score: without switching costs
Optimization Score: with switching costs
Point Forecast
+ Deterministic FHC
Probabilistic forecast
+ Stochastic FHC
Figure 5: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
(a) deterministic FHC algorithm
with point forecast
Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
$cost(FHC(v))$ 0.95 -0.94 -0.27
$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
(b) stochastic FHC algorithm
with probabilistic forecast
Correlation Coef. $ES$ $MASC_{V}$ $MASC_{H}$
$cost(FHC_{stoch.}(v))$ 0.63 -0.86 -0.29
$cost(FHC_{stoch.}(v))$ w/ switching costs 0.71 -0.06 0.47
Table 3: Correlation Coefficients for Forecast Metrics and Optimization Score
4 Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis in A.3, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions. In Figure 6, we can see that the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot. The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue). The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals. Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes. We assume that the oracle information given optimizing agent the confidence to charge at maximum capacity when the price and emissions are low. However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
The figure illustrates the difference in net load for two scenarios: one without switching costs and one with. When switching costs are considered, the optimal policy smooths the net load curve, resulting in a more stable profile with fewer fluctuations. The charging and discharging actions contribute to this smoothing, making the aggregate net load more stable. However, the presence of forecast errors exacerbates the switching costs. The aggregate EMS-controlled net load curve (in green) is not as smooth and flat as the optimal policy (in yellow). While it generally follows the optimal policy, it oscillates around it.
(a) Without switching costs
(b) With switching costs
Figure 6: The difference in net load between FHC with and without switching costs. Actual net load is shown in yellow. The base load is shown in green. The net load following the perfect forecast is shown in red. The electricity price is shown in blue.
Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction. This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run. In line with the theoretical analysis in Appendix A.3, the upper bound on the competitive difference of the FHC algorithm is dependent on the trade-off between switching costs and forecast errors.
While the MPC objective functions factors in the switching costs, we identify two factors that can lead to underperformance of FHC with low commitment period ($v_{O}=1$), particularly when probabilistic forecasts are used. Firstly, when the policy is computed for a finite period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the previously executed decision may be suboptimal given the updated information. A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable. Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier. The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances. However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
Other important observation is that the FHC algorithm with longer commitment periods only becomes viable when the optimization is computed for a set of scenarios, as we can see from the results in Figure 5. It is hypothesized, that the aggregation of policies over different scenarios provides a more robust policy that hedges against the forecast errors. Generally, within a forecast horizon, the nearest steps are predicted with higher accuracy than further steps. Longer commitment periods aggravate the impact of forecast errors due to time coupling, as the suboptimal decisions made in the past are amplified in the future. A stochastic FHC algorithm is better equipped to handle errors in the forecast, as it aggregates policies over different scenarios, therefore, the impact of forecast errors is mitigated and longer commitment periods are more viable.
5 Conclusion
To summarize, the inclusion of switching costs in the FHC algorithm introduces a trade-off between forecast errors and switching costs. While traditional FHC formulations can penalize switching costs in the objective function, longer commitment periods provide a natural mechanism to use a stable forecast and reduce the need for frequent adjustments. This approach is more viable when optimization is computed over a set of scenarios, since the aggregation of policies offers a more robust policy, hedging against increasing forecast errors in a forecast horizon.
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. Our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs. Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
Looking ahead, there are questions remaining to be addressed. For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy. Further research into the link between forecast quality and optimization performance is needed. Parameters such as forecast horizon, commitment period, and the number of scenarios are subject to a study for deeper understanding of their impact on the optimization performance and dependencies to the forecast accuracy and stability. It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading. From literature and current findings, methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors, however, they require more computational resources and are more complex to implement. An interesting avenue for future research would be defining the criteria for problems where these methods are most effective. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
Acknowledgments
This project has received funding from the European Union’s research and innovation programme Horizon Europe under the grant agreement No.101136211
Appendix A Cost Function and Optimization
A.1 Optimization score
Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization. The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. The optimization score is defined as:
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (10)
When grid-related KPIs are included, the average includes the grid score $D$. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score. The average score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
The normalized electricity cost, $C$, as delineated in Eq. 12, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq. 13 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
$C=\frac{c_{{submission}}}{c_{{nobattery}}}$ (12)
$c=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times T_{h}\right)$ (13)
Similarly, the normalized carbon emissions, $G$, is defined in Eq. 14 as the ratio of district carbon emissions for a given policy, $g_{{policy}}$, relative to the emissions in the aforementioned baseline scenario, $g_{{nobattery}}$. The emission metric $g$ is elaborated in Eq. 15 as the sum of carbon emissions, measured in $\left({kg}_{\mathrm{CO}_{2}\mathrm{e}}/\mathrm{kWh}\right)$, given by $E_{h}\times O_{h}$. Here, $O_{h}$ represents the carbon intensity for the hour $h$.
$G=\frac{g_{{submission}}}{g_{{nobattery}}}$ (14)
$g=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times O_{h}\right)$ (15)
Lastly, the evaluation metric includes a grid-related KPI, $D$. The metric follows grid-level objectives, such as the minimization of ramping and load factor. It is defined as the mean of the normalized ramping KPI, $R$, and the Load Factor KPI, $(1-L)$. The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq. 16.
$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (16)
The ramping KPI, $R$, reflects the smoothness of the district’s load profile. A low $R$ indicates a gradual increase in grid electricity demand even after self-generation becomes unavailable in the evening and early morning, while a high $R$ indicates abrupt changes in load on the grid, which may lead to unscheduled strain on grid infrastructure and potential blackouts due to supply deficits. It is calculated as the sum of the absolute difference of net electricity consumption between consecutive time steps:
$R=\sum_{t=1}^{8760}|E_{t}-E_{t-1}|$ (17)
The Load Factor, $L$, indicates the efficiency of electricity consumption and is bounded between 0 (very inefficient) and 1 (highly efficient). Thus, the goal is to minimize $(1-L)$. $L$ is calculated as the average ratio of monthly average to maximum net electricity consumption:
$L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (18)
A.2 Optimization Formulation
$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (19)
$=$ (20)
subject to
$SOC_{min}\leq SOC_{t}\leq SOC_{max},\quad\forall t,$ (21)
$x_{t}=x_{t}^{pos}+x_{t}^{neg},\quad\forall t,$ (22)
$-P_{max}\leq x_{t}^{neg}\leq 0,\quad\forall t,$ (23)
$SOC_{t}=SOC_{t-1}+$ (24)
$\eta_{charging}\cdot x_{t}^{pos}-\frac{1}{\eta_{discharging}}\cdot x_{t}^{neg},\quad\forall t.$ (25)
In this formulation:
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
    $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process. We assume that the weight factors are equal for all scenarios. In the deterministic case, the number of scenarios is equal to 1.
  • 3.
    $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively. Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
  • 4.
    $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
  • 5.
    $P_{max}$ is the maximum power with which the battery can be charged or discharged.
  • 6.
    $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
A.3 Theoretical Analysis of Performance Bounds
In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$. We focus on the cost difference between the optimal policy and the policy produced by the FHC algorithm. Specifically, we examine the impact of the commitment $v$, which defines how often the policy and forecast are updated. Our goal is to show how the performance bound is influenced by a trade-off between forecast accuracy and switching costs.
To illustrate this, we consider two models of forecast error: independent and identically distributed (i.i.d.) noise, noise with exponentially decaying correlation and a probabilistic forecast. We start with the deterministic case (point forecast) and then extend to the probabilistic case (set of scenarios). Following the derivation in Chen et al. [18], we derive the average-case performance of the FHC algorithm by taking the expectation of the cost difference with respect to the prediction noise $(e(t))^{T}_{t=1}$.
Here, OPT represents the optimal policy that would achieve the lowest possible cost if there were perfect information and no prediction errors.
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (26)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
$\mathbb{E}\operatorname{cost}(FHC)\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\left\|f_{v}\right\|^{\alpha}.$ (27)
The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition. The two noise models (i.i.d. and exponentially decaying correlation) are explained below.
Independent and Identically Distributed Noise
The i.i.d. noise is a simple and understandable baseline for forecast error, without time dependency. For i.i.d. noise, we have:
$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
The statistical properties of prediction errors in this context are:
$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
$\operatorname{trace}(R_{e})=\sigma^{2}$
This leads to the upper bound on the FHC cost:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\sigma^{\alpha}$ (28)
The upper bound on the competitive difference depends on $v$ only in the first term (in the denominator). Therefore, the optimal cost is minimized when $v=H$, i.e., the commitment level reaches the maximum allowable duration and corresponds to the forecast horizon length. The first term represents switching costs, so a shorter commitment period leads to more frequent updates and higher switching costs. The second term, representing i.i.d. noise, is independent of $v$.
Exponentially Decaying Correlation
A more realistic model of prediction noise uses correlations that decay exponentially over time. For this, given $a<1$, the Frobenius norm of the correlation structure is:
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (29)
We then express the prediction error covariance as:
$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\left\langle R_{e}^{1/2},f(s)\right\rangle^{2}$
$\leq\sum_{s=0}^{v}\left(\left\|R_{e}^{1/2}\right\|_{F}\|f(s)\|_{F}\right)^{2}$
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
Given that $h$ exhibits G-Lipschitz continuity ($\alpha=1$), the norm of the correlation function $\|f_{v}\|$ is bounded by:
$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (30)
Plugging this into the expression for the expected competitive difference of the FHC algorithm, we get:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (31)
$\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GT\sigma^{\alpha}\frac{c(1-a^{2(v+1)})/2}{1-a^{2}}$
$\leq\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GTc\sigma\left(\frac{a^{2v+2}-2}{2(a^{2}-1)}\right)$
To simplify, we group the parameters independent of $v$:
$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
Thus, the expected competitive difference of the FHC algorithm is:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{A}{v}+B(Ca^{2v}-1)$ (32)
The competitive difference behavior depends on two main components. The term $\frac{A}{v}$ decreases with $v$, while the term $B(Ca^{2v}-1)$ increases with $v$. Since $Ca^{2v}$ approaches 0 as $v$ increases, the second term is bound by $B$. Therefore, the minimal competitive difference is a trade-off between switching costs (first term) and forecast errors (second term).
Stochastic Optimization
For stochastic optimization with Fixed Horizon Control (FHC) applied across $n$ scenarios, we examine how the stochastic nature influences the competitive difference compared to an optimal solution. The expected cost for the stochastic algorithm across all scenarios is:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))=\sum_{i=1}^{n}\mathbb{E}\operatorname{cost}(FHC(v,i))$ (33)
Taking the expectation of the competitive difference across all scenarios:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (34)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\left[\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}\right]$
In the stochastic case, the forecast errors across the scenario ensemble would be lower in expectation than in any singular, deterministic scenario. This is because the stochastic approach averages out the errors across multiple scenarios, reducing the overall forecast error. Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
A.4 In-sample Performance
(a) Performance of the FHC algorithm with point forecast.
Study over combinations of forecast and optimization commitment between 1 and 12
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.486
2 0.489 0.489
3 0.491 0.492 0.491
4 0.492 0.492 0.494 0.492
5 0.493 0.494 0.494 0.494 0.493
6 0.493 0.493 0.493 0.495 0.495 0.493
7 0.493 0.494 0.495 0.495 0.495 0.495 0.493
8 0.493 0.493 0.495 0.493 0.495 0.496 0.495 0.493
9 0.494 0.495 0.494 0.495 0.495 0.495 0.495 0.496 0.494
10 0.494 0.494 0.495 0.495 0.494 0.495 0.496 0.495 0.496 0.494
11 0.494 0.495 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.494
12 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.495 0.496 0.496 0.495
(b) Performance of the FHC algorithm with stochastic forecast.
Study over combinations of forecast and optimization commitment between 1 and 12
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.798
2 0.805 0.800
3 0.809 0.806 0.803
4 0.809 0.805 0.807 0.802
5 0.811 0.807 0.806 0.805 0.805
6 0.812 0.807 0.805 0.805 0.806 0.805
7 0.810 0.808 0.807 0.805 0.806 0.808 0.805
8 0.808 0.806 0.808 0.802 0.807 0.807 0.807 0.804
9 0.811 0.809 0.806 0.805 0.806 0.807 0.808 0.806 0.805
10 0.811 0.807 0.807 0.805 0.806 0.807 0.808 0.806 0.808 0.806
11 0.810 0.808 0.808 0.805 0.807 0.808 0.807 0.807 0.808 0.808 0.806
12 0.813 0.809 0.807 0.805 0.806 0.808 0.807 0.806 0.807 0.808 0.808 0.807
Table 4: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
Accuracy
Optimization Score: without switching costs
Stability: vertical (top) and horizontal (bottom)
Optimization Score: with switching costs
Point Forecast + Deterministic FHC
Probabilistic forecast + Stochastic FHC
Figure 7: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from in-sample buildings