### v5

# Balancing Forecast Accuracy and Switching Costs in Online Optimization of Energy Management Systems

Abstract
This study investigates the integration of forecasting and optimization in energy management systems, with a focus on the role of switching costs, defined as penalties incurred from frequent operational adjustments. We develop a theoretical and empirical framework to examine how forecast accuracy and stability interact with switching costs in online decision-making settings. Our analysis spans both deterministic and stochastic optimization approaches, using point and probabilistic forecasts. A novel metric for measuring temporal consistency in probabilistic forecasts is introduced, and the framework is validated in a real-world battery scheduling case based on the CityLearn 2022 challenge. Results show that switching costs significantly alter the trade-off between forecast accuracy and stability, and that more stable forecasts can reduce the performance loss due to switching. Contrary to common practice, the findings suggest that, under non-negligible switching costs, longer commitment periods may lead to better overall outcomes. These insights have practical implications for the design of intelligent, forecast-aware energy management systems.
keywords:
Energy Management Systems, Forecasting, Optimization, Switching Costs, Forecast Stability, Model Predictive Control
[…]
[cor1]Corresponding author
1 Introduction
Managing energy assets within a grid system presents a challenging task characterized by decision-making under uncertainty. The inherent dynamics and stochasticity of the environment make this a complex system, constantly evolving and demanding decisions to be made with incomplete knowledge of the future over limited time windows, a scenario typically addressed through online optimization.
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging Kriekinge et al. [2021], Tungom et al. [2024], battery scheduling Ruddick et al. [2024], Liu et al. [2023], and energy and flexibility dispatch problems Qu et al. [2024], Liu et al. [2025]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down Wu et al. [2013], managing server load balancing Gupta [2021], Lin et al. [2012], or covering network fees associated with energy trading Yang et al. [2023], Levorato et al. [2022].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information.
A key consideration in these systems is determining the optimal commitment period, meaning how long one can commit to a given policy before recalculating based on updated forecasts. This commitment level directly influences system performance, especially when switching costs are present. Additionally, forecast stability, the consistency of predictions across successive updates, plays a critical role in the reliability and efficacy of the optimization process. Striking a balance between forecast accuracy, stability, and the impact of switching costs is crucial. This paper explores how this balance shapes system design and informs operational strategies. Consequently, this research aims to answer the following questions:
  • Does presence of switching costs affect the optimal commitment level of the policy?
  • How do forecast accuracy and stability influence the downstream performance of energy management systems?
  • How can integrated systems of forecasting and optimization be designed to effectively balance these considerations for improved decision-making in energy management?
1.1 Related work
Literature solutions for integrating forecasting and optimization are classified based on the level of integration between the two components. This classification includes three categories: Direct, Indirect, and Semi-direct methods. Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’. Such approach is proposed in Elmachtoub and Grigas [2022], with optimization cost integrated directly into the forecast loss function to align forecasts with end-use optimization. This optimizes the prediction model in the forecasting stage for better decision-making in the optimization stage. The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges. There are other works that design an alternative task-specific loss function, such as reported in Mandi et al. [2020], and tested on a real-world energy management problem.
Indirect methods regard the two sub-problems as separate, also referred in literature as ‘predict, then optimize’. This approach is regarded as standard and is widely applied in practice. In the study by Vanderschueren et al. [2022], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification. Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of a classification algorithm outperforms the integrated approach. Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside the training process. An example is found in the work by Kazmi and Paskevich [2023], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is Huang et al. [2025], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task. Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment. A relevant study by Prat et al. [2024] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach. The study introduces a verifiable condition to check if a selected planning horizon is sufficiently long. Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The frequency of retraining is studied by Spiliotis and Petropoulos [2024], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM Ke et al. [2017]).
We are particularly interested in studying the model deployment frequency. More specifically, we look at the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization. In literature, the FHC/RHC algorithm is often referred to as Model Predictive Control (MPC), which is a well-established method for solving online multi-step optimization problems in control theory.
(a) FHC
(b) RHC
Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control
The challenge of noisy predictions in Online Convex Optimization (OCO) has been extensively studied, notably by Chen et al. [2015], who emphasized the difficulty of designing robust online algorithms under uncertainty. In response, various strategies have emerged to mitigate the impact of prediction noise. For instance, Averaging Fixed Horizon Control (AFHC) and its generalized form, Committed Horizon Control (CHC) Chen et al. [2016], employ aggregation across multiple plan revisions to achieve sublinear regret and greater robustness. Likewise, the Feasible Fixed Horizon Control (FFHC) algorithm Christianson et al. [2022] introduces regularization to ensure feasibility in multi-interval settings while addressing ramping-related switching costs. In this study, we adopt the classical Fixed Horizon Control (FHC) algorithm with a fixed commitment level as the foundational optimization scheme. FHC is not only conceptually simple but also widely adopted in energy management systems, particularly in implementations of Model Predictive Control (MPC) Ahmad et al. [2023]. While previous works have primarily focused on aggregating or averaging policies to improve performance, our focus shifts toward examining how the frequency of re-optimization, i.e., the commitment level, interacts with forecast stability and switching costs.
Forecast stability itself has recently gained attention as a key dimension in sequential decision-making. Building on the taxonomy proposed by Godahewa et al. [2023], which distinguishes between vertical stability (variation across forecast updates for the same time step) and horizontal stability (variation within the forecast horizon at a single time step), we explore how these notions influence system performance in the energy domain. Though similar concerns have been addressed in supply chain literature, where freezing intervals Sahin et al. [2013], Xie et al. [2003] and smoothing techniques Ho* [2005] aim to reduce system nervousness and the bullwhip effect Wang and Disney [2016], energy systems impose different requirements. Forecasts must be updated more frequently and switching costs (such as ramping costs Wu et al. [2013], Tanaka [2006] or toggling server states Lin et al. [2012], Lee et al. [2021]) are often tied to physical infrastructure.
Despite the relevance of these factors, few studies have systematically explored the link between commitment strategies, forecast stability, and switching costs in energy systems. This study addresses that gap by investigating how varying commitment levels affect performance, particularly in energy storage scheduling and real-time optimization contexts. By doing so, we aim to inform the design of more robust and responsive decision-making frameworks in energy management.
1.2 Contents and Contributions
The principle contributions are as follows:
  • First, we formally define the problem and the framework for online optimization with switching costs, introducing the Fixed Horizon Control (FHC) algorithm and the concept of forecast stability. We develop the theoretical analysis that highlights how the commitment level influences the trade-off between forecast accuracy and switching costs.
  • We propose a novel metric called Scenario Distribution Change (SDC) for evaluating the stability of probabilistic scenario sets, extending the concept of forecast stability from point forecasts to probabilistic forecasts. This metric allows for measuring both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates).
  • We conduct an empirical evaluation using the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Through extensive testing of different commitment periods, we analyze the relationship between forecast accuracy, stability, and policy performance, demonstrating that in the presence of switching costs, longer periods of commitment can lead to improved decision-making and system performance.
  • Lastly, we discuss the implications of our findings for the design of energy management systems, highlighting how the balance between forecast accuracy, stability, and switching costs shapes optimal operational strategies in real-world applications.
2 Problem Formulation and Framework
This section begins by introducing a general problem framework for online optimization with switching costs, defining the mathematical formulation and related key concepts. Next, we present the Fixed Horizon Control algorithm as the primary solution method, discussing how commitment periods affect decision-making. The concept of forecast stability is also introduced to analyze its impact on system performance.
2.1 Online Optimization with Switching Costs
Online optimization with switching costs is a class of problems characterized by sequential decision-making under uncertainty. In these problems, decisions must be made in real-time as new information becomes available, with only partial knowledge of future conditions. What makes this class of problems particularly challenging is the presence of time-coupling effects, where decisions at one time step directly affect the state of the system and the options available at subsequent time steps. The general problem is defined as finding an optimal state-dependent policy $X_{t}(S_{t}|\theta^{LA})$. The decision-maker faces a sequence of decisions, where each decision is based on the current system state $S_{t}$ and the direct look-ahead approximation $\theta^{LA}$. Using the framework proposed by Powell [2019], the problem can be formulated as follows:
$X_{t}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
$\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
$+\beta\left\lVert\tilde{S}_{tt^{\prime}}-\tilde{S}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
In this formulation:
  • $\mathbb{C}(\textbf{x};\textbf{S})$ denotes the total cost function, which is a function of the decision variables x and the system state S
  • $W_{i}$ represents the weight associated with different cost components or scenarios. For deterministic optimization, only one scenario exists, while stochastic approaches consider multiple weighted scenarios.
  • $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ captures the direct operational cost at time $t^{\prime}$ when following decision $x_{tt^{\prime}}$ under forecast scenario $\theta^{LA}_{i,tt^{\prime}}$.
[…]
2.2 Fixed Horizon Control Algorithm
To address online optimization problems with switching costs, the Fixed Horizon Control (FHC) algorithm is employed. FHC is a variant of model predictive control with a fixed commitment horizon. At each decision point, the algorithm computes an optimal policy for the current state of the system and forecasts for the next $H$ time steps, determining the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$.
A key parameter in FHC is the commitment period $v$, which defines how many steps of the computed policy are implemented before recomputing. Every $v$ time steps, the optimization is rerun to generate a new plan. When $v=1$, the algorithm becomes Receding Horizon Control (RHC), where optimization is performed at every time step, implementing only the first step of each plan. Both the forecast and optimization follow a rolling origin setup and require updates at certain intervals while committing to $v$ steps. At each time step, the decision-maker faces a choice to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. This decision framework is illustrated in Table 1. It is important to note that not all quadrants in this decision matrix are equally practical or beneficial. Specifically, updating the forecast without updating the optimization plan (bottom-left quadrant) offers limited value, as new information is obtained but not acted upon. Conversely, updating the optimization without updating the forecast (top-right quadrant) can be valuable in situations where the system state changes in ways unrelated to the forecast variables. In most practical implementations, the most logical configurations are either reusing both forecast and plan (top-left) or revising both simultaneously (bottom-right).
Table 1: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
Optimization Decision
Forecasting Decision Reuse Revise
Reuse Continue with current plan Evaluate
& adjust current plan
Revise Evaluate &
retain current plan
Re-evaluate both
forecasting & plan
2.2.1 Rolling Origin Window
In practical applications, forecasts are issued periodically for a finite number of future time steps. This process runs iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps. Figure 2 illustrates this approach, which involves continuously moving the time window forward by a certain step size after each forecasting iteration.
  Input Data
 $\rightarrow$ Forecast Horizon
$t_{0}$$t_{1}$$t_{2}$$t_{3}$$t_{4}$$t_{5}$$t_{6}$$t_{7}$$t_{8}$orecast1Forecast2Forecast3Forecast4Forecast Origins
Figure 2: Rolling Origin Forecasting with Fixed Input Window and Forecast Horizon. The diagram illustrates how forecasts with horizon $H=2$ are updated every $v_{f}=1$ time steps, creating overlaps between prediction windows.
The rolling horizon setup enables prediction updates using the most recent information. When the commitment period is shorter than the forecast horizon, regular updates create overlaps between prediction windows, as forecasts for a particular future time point are released in multiple revisions. These overlaps are considered when evaluating forecast stability.
2.2.2 Stability
Forecast stability is defined as the variability of predictions over time. Following the categorization proposed by Godahewa et al. [2023], we distinguish between vertical and horizontal stability, as illustrated in Figure 3.
(a) Vertical Stability
(b) Horizontal Stability
Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions.
Vertical stability refers to the consistency between predictions made at different origins for the same target time period. Horizontal stability, on the other hand, characterizes the variance of predictions within the same forecast window.
3 Theoretical analysis of performance bounds
This section presents an analysis of the performance bounds for Fixed Horizon Control (FHC) algorithms with commitment level $v$. We derive upper bounds on the competitive difference, the gap between FHC and the optimal clairvoyant policy. Our analysis reveals a fundamental trade-off between forecast accuracy and switching costs that depends critically on how frequently policies are updated. We first develop these bounds in a deterministic setting, then extend them to stochastic optimization where forecasts are represented as scenario ensembles. Finally, we establish the theoretical connection between forecast stability and switching costs, demonstrating why stable forecasts can significantly improve performance in environments with high switching costs.
We adopt the framework introduced by Chen et al. [2016] for analyzing online convex optimization (OCO) problems with switching costs and noisy predictions. In this setting, a decision-maker sequentially selects actions $x_{t}\in\mathcal{F}\subset\mathbb{R}^{n}$ over a time horizon $T$, facing unknown convex cost functions $h_{t}(x_{t})$ and incurring a switching cost $\beta\|S_{t}-S_{t-1}\|$.
$\operatorname{cost}(ALG)=\sum_{t=1}^{T}h_{t}(x_{t})+\sum_{t=1}^{T}\beta\|S_{t}-S_{t-1}\|$ (2)
The true state $y_{t}$ at each time $t$ is unknown at the time of decision-making and must be predicted. Forecasts $\hat{y}_{t|\tau}$ are generated at time $\tau<t$ and are corrupted by noise, modeled as:
$y_{t}-\hat{y}_{t|\tau}=\sum_{s=\tau+1}^{t}f(t-s)e(s),$ (3)
where $e(s)$ are i.i.d. zero-mean noise terms with covariance matrix $R_{e}$, and $f(\cdot)$ is a deterministic impulse response function capturing the correlation structure of the forecast errors.
The performance of an online algorithm is assessed using the competitive difference, defined as the expected cost gap between the online algorithm and the offline optimal (oracle) policy. Formally, an online algorithm ALG, is said to have a competitive difference of $\rho(T)$ if:
$\sup_{\hat{y}}\mathbb{E}_{e}\left[\operatorname{cost}(\text{ALG})-\operatorname{cost}(\text{OPT})\right]\leq\rho(T)$ (4)
where the expectation is taken with respect to the prediction noise sequence $\{e(t)\}_{t=1}^{T}$. The offline optimal policy OPT has full access to future outcomes and is thus not affected by forecast errors. In contrast, the performance of online algorithms like Fixed Horizon Control (FHC) or Receding Horizon Control (RHC) is directly impacted by the structure and magnitude of the forecast noise.
Following this framework, we derive bounds on the expected performance degradation of FHC-type algorithms as a function of the commitment level $v$, the forecast error correlation $\|f_{v}\|$, and the switching cost penalty $\beta$. As a starting point, we refer to Theorem 1 in Chen et al. [2016], and adapt it to our context.
$\operatorname{cost}(FHC(v))\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\|S_{t}-S_{t\mid\phi_{k}(t)}\|$ (5)
where $M$ is the number of updates, $k$ is the commitment level, $D$ is the maximum switching cost, and $G$ is the Lipschitz constant of the cost function
Here, in contrast to the original paper, we derive the expectation over the forecast horizon for the FHC algorithm with commitment level $v$:
$\mathbb{E}[\operatorname{cost}(FHC(v))]$ $\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\mathbb{E}\left[\|S_{t}-S_{t\mid t-k}\|\right]$ (6)
$\leq\operatorname{cost}(OPT)+2(T/v)\beta D+2GT\cdot\|f_{v}\|^{\alpha}$ (7)
as $M_{k}$ is the number of optimization updates, which is equal to $T/v$ for the FHC(v) algorithm.
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (8)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|S_{t}-S_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
$\mathbb{E}\operatorname{cost}(FHC)\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\left\|f_{v}\right\|^{\alpha}.$ (9)
The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition. We assume $\alpha\geq 1$ to ensure that the cost function is convex.
3.1 Deterministic Optimization
In practical applications, forecast errors often exhibit exponentially decaying temporal correlations. We model this behavior using a decay parameter $a\in(0,1)$ and define the Frobenius norm of the correlation impulse response as:
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (10)
where $c$ is a constant scaling factor and $s$ is the time lag between the forecast target time and forecast issuance. We assume zero-mean i.i.d. noise vectors $e(s)$ with covariance $\mathbb{E}[e(s)e(s)^{T}]=R_{e}$ and trace $\operatorname{trace}(R_{e})=\sigma^{2}$. The cumulative forecast error norm up to commitment level $v$ becomes:
$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\|R_{e}^{1/2}\|_{F}^{2}\cdot\|f(s)\|_{F}^{2}$
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\cdot\frac{1-a^{2(v+1)}}{1-a^{2}}$
Taking the square root and using the Lipschitz continuity of $h$ with constant $G$, we have:
$\|f_{v}\|=c\sigma\sqrt{\frac{1-a^{2(v+1)}}{1-a^{2}}}$ (11)
Substituting into the expected cost bound for the FHC algorithm, we obtain:
$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{2T\beta D}{v}+2GT\|f_{v}\|$ (12)
Define:
$A$ $=2T\beta D$
$B$ $=\frac{2GTc\sigma}{\sqrt{1-a^{2}}}$
Then the bound becomes:
$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{A}{v}+B\cdot\sqrt{1-a^{2(v+1)}}$ (13)
This formulation clearly reflects the trade-off. The term $\frac{A}{v}$ is the cost of policy switching and decreases with larger commitment $v$. The term $B\cdot\sqrt{1-a^{2(v+1)}}$ captures forecast error accumulation and increases with $v$ since $a^{2(v+1)}\to 0$ as $v$ increases. The role of each parameter is summarized in Table 2. Therefore, the expected competitive difference exhibits a U-shaped curve with respect to $v$, where a minimal cost is achieved by balancing the cost of frequent switching and the degradation due to forecast staleness. This quantifies the intuition that updating too frequently incurs high switching costs, while updating too infrequently leads to performance loss due to outdated forecasts. This function run with different sets of parameters is demonstrated in Figure 7. It is observed that the low switching costs lead to a linear increase in the expected competitive difference, while the high switching costs lead to a formation of a U-shaped curve.
Table 2: Key parameters in performance bound
Symbol Represents Interpretation
$A$ Switching Cost Weight Scales with horizon $T$, penalty $\beta$, and max switching distance $D$. Larger $A$ increases the switching cost.
$B$ Forecast Error Sensitivity Incorporates cost weight $G$, horizon $T$, and noise parameters. Amplifies penalty for forecast inaccuracy.
$a$ Correlation Decay Rate Controls how quickly forecast errors become uncorrelated over time. Lower values of $a$ mean errors at future steps are largely independent.
Figure 4: Illustration of the FHC expected competitive difference as a function of the commitment level for (a) Low and High switching costs with predictions of exponentially decaying correlations, and (b) Stochastic and Deterministic optimizations. Stochastic scenarios exhibit a lower level of forecast error sensitivity ($B_{s}<B$), shifting the optimal commitment level to higher values.
3.2 Stochastic Optimization over Scenarios
We now analyze the performance of Fixed Horizon Control (FHC) under a stochastic optimization setting where forecasts are provided as a set of $n$ equiprobable scenarios, denoted $\{\xi_{i}\}_{i=1}^{n}$, each representing a possible trajectory of future values over the horizon $v$. For each scenario $\xi_{i}$, the FHC algorithm yields a policy $x^{(i)}=(x^{(i)}_{t})_{t=1}^{T}$. The overall cost is computed by averaging over the scenario outcomes:
$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\operatorname{cost}(FHC(v,\xi_{i}))\right].$ (14)
Let $f_{v}^{(i)}\in\mathbb{R}^{d}$ denote the cumulative forecast error over horizon $v$ under scenario $\xi_{i}$. The expected cost includes a term of the form $\|f_{v}^{(i)}\|^{\alpha}$, which contributes to the second term in the performance bound.
Assuming $\alpha\geq 1$, the function $x\mapsto\|x\|^{\alpha}$ is convex, and Jensen’s inequality yields:
$\left\|\frac{1}{n}\sum_{i=1}^{n}f_{v}^{(i)}\right\|^{\alpha}\leq\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (15)
This inequality implies that the effective contribution of forecast uncertainty to the cost is lower in the stochastic case than in the deterministic case, since:
$2GT\left\|\mathbb{E}[f_{v}^{(i)}]\right\|^{\alpha}\leq 2GT\cdot\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (16)
Thus, scenario averaging introduces a variance reduction effect, which improves robustness to forecast noise. This insight leads to the following bound:
$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]\leq\mathbb{E}\left[\operatorname{cost}(OPT)\right]+\frac{A}{v}+B_{s}(1-a^{2(v+1)})^{1/2},$ (17)
where $B_{s}<B$ captures the effect of scenario averaging. This result highlights a key theoretical benefit of stochastic FHC: the aggregation over multiple forecast paths leads to a lower effective forecast variance and therefore lower forecast error sensitivity. It is shown in Figure 4(b), that the vertex of the U-shaped curve shifts to higher values of $v$ in the stochastic case. The fundamental trade-off between switching costs (decreasing with $v$) and forecast accuracy (typically decreasing with more frequent updates) persists, but the stochastic case enables a longer commitment horizon before the forecast error term dominates.
3.3 Role of Forecast Stability in Optimization Performance
The theoretical performance bounds in Equation (9) consist of two main components: (i) a term proportional to the switching cost, $\frac{2T\beta D}{v}$, and (ii) a forecast error penalty, $2GT\|f_{v}\|^{\alpha}$. While the second term captures the accuracy of the forecast over the horizon $v$, the first term reflects the cost incurred by changing policies across planning updates. In this section, we formalize the connection between the stability of forecasts and the incurred switching costs. Let $x_{t}$ denote the decision policy applied at time $t$, and let this policy be a deterministic function of the forecast $\hat{y}_{t}$, such that:
$x_{t}=\mathcal{M}(\hat{y}_{t}),$ (18)
where $\mathcal{M}:\mathbb{R}^{H}\to\mathbb{R}^{d}$ is a policy generation function (e.g., the solution to an optimization problem over horizon $H$), and is assumed to be Lipschitz continuous with constant $L_{\mathcal{M}}$:
$\|\mathcal{M}(\hat{y}_{t})-\mathcal{M}(\hat{y}_{t-v})\|\leq L_{\mathcal{M}}\cdot\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (19)
Using this, the switching cost term can be bounded as:
$\sum_{t=1}^{T/v}\beta\|S_{t}-S_{t-v}\|\leq\beta L_{\mathcal{M}}\sum_{t=1}^{T/v}\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (20)
The term $\|\hat{y}_{t}-\hat{y}_{t-v}\|$ represents the change in forecast between updates, which corresponds to the concept of vertical forecast stability—the consistency of predictions made at different time origins for the same target time. A common empirical proxy for this is the Mean Absolute Change (MAC) metric:
$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}|\hat{y}_{t+i|t}-\hat{y}_{t+i|t-v}|.$ (21)
Thus, improving the vertical stability of forecasts directly reduces the magnitude of policy changes, and therefore the incurred switching costs. This establishes a theoretical link between empirical measures of forecast stability and the first term in the performance bound (9). The MAC metric and appropriate stability metrics for the stochastic case are discussed in detail in the next section.
4 Experimentation and Results
In this section, we transition from theoretical analysis to practical implementation and evaluation. While the theoretical model provides important insights, real-world energy management systems face additional complexities including non-trivial forecast error correlations, potential non-convexity in optimization problems, and the challenges of coordinating multiple distributed assets. We present empirical results from applying the proposed framework to a battery scheduling problem, evaluating how different commitment periods affect system performance under varying forecast qualities. First, the current section outlines the experimental setup and case study, followed by a detailed analysis of the results. The analysis focuses on the relationship between forecast accuracy, stability, and the impact of switching costs on the optimal decisions made in the battery scheduling process.
4.1 Application to Battery Scheduling in Energy Management Systems
The decision-maker must balance the benefits of plan revisions against the costs of switching. This balance is particularly critical in energy management systems, where rapid changes in operation can strain infrastructure, reduce equipment lifespan, and lead to additional expenses or inefficiencies. To investigate these trade-offs in a realistic setting, we apply the concepts of online optimization with switching costs to battery storage management within a multi-agent energy management system. Our case study is based on the CityLearn Challenge 2022 Nweye et al. [2022], which provides a comprehensive dataset from an actual grid-interactive community.
The central challenge involves multi-agent scheduling of energy storage across a one-year period with hourly granularity. Figure 5 illustrates the problem setup. The data of 17 buildings used in this study is derived from a real-world zero net energy community in Fontana, California, USA, previously studied for grid integration of zero net energy developments as part of the California Solar Initiative program Narayanamurthy et al. [2016]. Each building has a battery with specific charging/discharging characteristics, and the system must coordinate these resources efficiently. The objective is to develop control policies that minimize grid electricity costs, carbon emissions and ramping while increasing the load factor. The associated ramping costs constitute the switching costs in this context. The combined objective score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
where $C$ are electricity costs, $G$ are carbon emissions, and $D$ is the grid score. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. Full details of the objective score and relevant constraints are provided in A.1.
The proposed framework was implemented using the CityLearn simulation environment, an open-source platform designed for evaluating energy management systems in grid-interactive communities [^1] https://github.com/intelligent-environments-lab/CityLearn. Originally created for reinforcement learning (RL), the platform now also supports alternative methods like Model Predictive Control (MPC) and rule-based control. It simulates a range of energy assets, including batteries, solar panels, heat pumps, and electric vehicles, all of which play a role in optimizing energy consumption across a grid-connected community.
Figure 5: System setup (a) and data flow (b) for the battery scheduling problem. The data flow diagram illustrates how input features are processed through forecasting models with varying update frequencies to generate load predictions, which then inform the optimization stage to produce battery control actions.
4.2 Implementations of Forecasting and Optimization
The system architecture consists of two primary components: forecasting and optimization, as shown in Figure 5(b). The forecasting module generates load predictions, which serve as inputs to the optimization model. Both forecasting and optimization follow a rolling horizon approach. While the commitment period $v$ is commonly the same for both stages, updates are not inherently required to be synchronized. Therefore, we distinguish between $v_{F}$ (forecast update frequency) and $v_{O}$ (optimization update frequency) when necessary.
For the forecasting task, we predict the aggregate uncontrollable load of the energy community. The training set comprises data of 5 buildings, with the remaining 12 buildings used for validation and testing. The forecasts are generated using a combination of gradient boosting regression trees (LightGBM Ke et al. [2017]) and linear least squares regression. The method follows the procedure of the winning solution, outlined in Nweye et al. [2022]. We train the forecasting model across all buildings, then apply it to 7 out-of-sample buildings for testing, a practice commonly used in scenarios with limited historical data but similar building characteristics, as discussed in Grabner et al. [2023]. The forecasts are generated for a 24-hour horizon, with 24 separate models predicting each hour. To account for forecast uncertainty, we generate multiple forecast scenarios by adding Gaussian noise to the base predictions.
The optimization model, implemented with Pyomo Hart et al. [2017] and solved using Gurobi Gurobi Optimization, LLC [2024], aims to minimize a composite cost function that incorporates electricity costs, ramping penalties, switching costs, and carbon emissions. The optimization problem is solved for the same 24-hour horizon as the forecasting model, subject to revision every $v_{O}$ steps. At the optimization stage, both deterministic and stochastic forecast settings are considered. In the deterministic case, a single forecast trajectory is used to determine the optimal control policy. In the stochastic case, multiple forecast scenarios are considered, and the optimization is performed over 75 scenarios, minimizing the expected cost. This allows the system to better handle forecast uncertainty and improves robustness by optimizing expected performance across scenarios.
The optimization and forecasting steps are coordinated through the CityLearn platform, which integrates data from multiple sources, including weather forecasts, building characteristics, and historical consumption patterns. The resulting control policies, which dictate how the batteries should charge and discharge, are evaluated based on their ability to reduce grid electricity costs, minimize ramping costs, and lower carbon emissions. The implementation is available on GitHub [^2] https://github.com/ujohn33/Predict-Optimize-Revise and is fully reproducible, with optimization logs available upon request. The detailed mathematical formulation of the optimization problem and the constraints imposed on the system are provided in A.2.
4.3 Evaluation metrics
Similarly to the evaluation in [Van Belle et al., 2023], a stability metric is measured for point forecasts using the mean absolute change (MAC). The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the variation of predictions within the horizon window — ${MAC}_{H}$
$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|$ (23)
$\operatorname{MAC}_{H}=\frac{1}{H-1}\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|$ (24)
where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
4.3.1 Stability of Probabilistic Forecasts
In stochastic optimization, scenario stability measures the consistency of probabilistic forecast updates over time. We propose the Scenario Distribution Change (SDC) metric, adapting the Wasserstein Distance, also referred to as Earth Mover Distance, to assess the stability of scenario sets. SDC quantifies the average change between successive scenario updates, providing a clear measure of stability for probabilistic forecasts in dynamic systems.
The Wasserstein distance formulation is symmetric, adheres to the triangle inequality, and effectively compares probabilistic scenario sets. Unlike metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon distance, SDC holds for non-overlapping distributions, making it particularly relevant for applications requiring forecast stability. By incorporating the dimension of the metric space, SDC is conceptually analogous to the Mean Absolute Change (MAC) used for point forecasts.
To assess stability across time, SDC evaluates both vertical stability (across forecast updates) and horizontal stability (within a forecast horizon). For $N$ number of scenarios and a forecasting horizon $H$, these are defined as:
Vertical Stability (SDCV):
$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (25)
Horizontal Stability (SDCH):
$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (26)
4.3.2 Accuracy of Point Forecasts
For point forecasts, the MAE is utilized to quantify the forecast accuracy. It calculates the average absolute difference between each forecasted value and the corresponding actual value. Mathematically, it is expressed as:
$\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (27)
4.3.3 Accuracy of Probabilistic Forecasts
For probabilistic forecasts, we employ the Energy Score (ES). The ES is a multivariate generalization of the continuous ranked probability score (CRPS), a widely used metric for evaluating probabilistic forecasts. The ES measures the distance between the forecasted distribution and the actual value. The lower the ES, the better the forecast. The ES is defined as:
$ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (28)
4.4 Results
According to the decision matrix in Table 1, at every time step the forecast and optimization plans can be updated or retained. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment, $v_{F}$ and $v_{O}$, respectively. The result is shown in Table 3. The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
Table 3: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours
Optimization Commitment Period ($v_{O}$)
Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
Panel A: Deterministic FHC with point forecast
1 0.899
2 0.910 0.904
3 0.911 0.907 0.904
4 0.912 0.908 0.907 0.905
5 0.913 0.910 0.908 0.908 0.903
6 0.912 0.910 0.907 0.908 0.908 0.906
7 0.912 0.910 0.907 0.910 0.908 0.909 0.902
8 0.910 0.907 0.908 0.905 0.906 0.908 0.902 0.903
9 0.911 0.909 0.906 0.908 0.908 0.908 0.904 0.908 0.903
10 0.913 0.909 0.909 0.907 0.906 0.908 0.903 0.906 0.905 0.905
11 0.912 0.910 0.909 0.910 0.908 0.909 0.905 0.906 0.906 0.907 0.903
12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
Panel B: Stochastic FHC with probabilistic forecast
1 0.875
2 0.888 0.875
3 0.894 0.886 0.873
4 0.895 0.885 0.882 0.874
5 0.899 0.892 0.885 0.886 0.875
6 0.901 0.891 0.881 0.883 0.884 0.876
7 0.901 0.896 0.887 0.888 0.886 0.883 0.875
8 0.900 0.894 0.887 0.884 0.884 0.885 0.882 0.874
9 0.902 0.894 0.883 0.887 0.884 0.882 0.883 0.880 0.876
10 0.903 0.894 0.888 0.887 0.881 0.884 0.883 0.880 0.883 0.876
11 0.902 0.895 0.889 0.887 0.885 0.885 0.884 0.880 0.884 0.882 0.876
12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
Note: Diagonal elements represent equal commitment periods for both forecast and optimization. Lower values indicate better performance. Empty cells (—) represent invalid combinations where $v_{O}>v_{F}$.
Accuracy
Stability: vertical (top) and horizontal (bottom)
Optimization Score: without switching costs
Optimization Score: with switching costs
Point Forecast
+ Deterministic FHC
Probabilistic forecast
+ Stochastic FHC
Figure 6: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 6, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.3. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours. For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in Nweye et al. [2022]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation. Nevertheless, as seen in the optimization score plots, the global minimum is reached at a shorter commitment period, indicating that updating the schedule more frequently is not only beneficial for using the most accurate forecast but also for correcting the previous decisions.
Furthermore, it is observed that the presence of switching costs has a significant impact on the relation of performance to the commitment period. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. Similar findings are devised from the average-case analysis in Section 3.1, where higher switching costs can potentially lead to a formation of a minimum beyond lowest commitment period. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period. This observation is notable as, conventionally, the best performance is achieved with the shortest revision periods. This observation is also consistent with the theoretical analysis in Section 3.2, as the stochastic optimization reduces sensitivity to forecast errors, leading to a more stable performance.
These observations are supported by the correlation coefficients in the Table 4 with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
Figure 7 demonstrates an example from the Citylearn case when the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot. The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue). The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals. Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes. We assume that the oracle information gives the optimizing agent the confidence to charge at maximum capacity when the price and emissions are low. However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
The figure illustrates the difference in net load for two scenarios: one without switching costs and one with. When switching costs are considered, the optimal policy smooths the net load curve, resulting in a more stable profile with fewer fluctuations. The charging and discharging actions contribute to this smoothing, making the aggregate net load more stable. However, the presence of forecast errors exacerbates the switching costs. The aggregate EMS-controlled net load curve (in green) is not as smooth and flat as the optimal policy (in yellow). While it generally follows the optimal policy, it oscillates around it.
Table 4: Correlation Coefficients between Forecast Metrics and Optimization Performance
Point Forecast (Deterministic FHC) Probabilistic Forecast (Stochastic FHC)
Correlation Coefficient $MAE$ $MAC_{V}$ $MAC_{H}$ $ES$ $SDC_{V}$ $SDC_{H}$
Without switching costs 0.95 -0.94 -0.27 0.63 -0.86 -0.29
With switching costs 0.70 -0.62 0.17 0.71 -0.06 0.47
(a) Without switching costs
(b) With switching costs
Figure 7: Net load comparison: Without switching costs, energy management with perfect forecast (yellow) generates high peaks versus conservative forecast-based control (blue). With switching costs, the policy smoothes fluctuations. The base load, being the prediction target variable, is shown in blue. Price and emissions signals are shown in the bottom subplot.
5 Discussion
This study provides theoretical and empirical evidence for the nuanced role that commitment periods play in forecast-based optimization under switching costs. Our results validate the theoretical performance bounds developed in Section 3, showing that the trade-off between forecast inaccuracy and switching costs emerges most prominently when switching penalties are non-negligible. Notably, this trade-off is not universal: in problems without meaningful switching costs, shorter commitment periods are almost always preferred. However, when switching costs are significant, such as in systems with physical constraints, ramping penalties, or contract-driven operational inertia, the choice of commitment level becomes a key decision variable in its own right.
The presence of switching costs changes the optimization landscape in fundamental ways. It introduces temporal coupling between successive decisions, requiring current actions to anticipate the cost of future revisions. This often leads to smoother and more conservative policies, as illustrated in Figure 7, the addition of switching costs results in a visibly flattened net load curve. Such behavior is desirable in many real-world settings: battery degradation, generator ramping, thermal comfort zones, and load shifting incentives all impose implicit or explicit switching penalties that must be carefully managed. Although such costs are context-specific and often application-dependent, they are far from rare in modern energy systems.
Importantly, we show that this trade-off is more tractable under stochastic optimization. Section 3.2 demonstrates that scenario averaging introduces a form of regularization, lowering the effective sensitivity to forecast errors. This effect is reflected in both the theoretical bounds and the experimental results. In practice, this means that higher commitment levels, normally problematic due to the accumulation of outdated forecast errors, can be made viable when robust scenario-based policies are used. The performance gains from longer commitment periods become visible only when switching costs are high and stochastic optimization is employed, as shown in Figure 6. This aligns with the theoretical prediction that the second term in the cost bound (forecast error contribution) becomes less dominant when multiple forecast scenarios are aggregated.
From an operational perspective, this has several implications. First, minimizing switching costs is not solely about reducing the number of policy changes. It also requires designing forecasts that are stable across time. As formalized in Section 3.3, vertical forecast stability directly mitigates the switching cost penalty. Stability metrics such as Mean Absolute Change (MAC) or Scenario Distribution Change (SDC) can thus serve as practical proxies to monitor and optimize for reduced control variability in real systems. Second, while the importance of forecast accuracy has long been emphasized in predictive control, this study shows that stability is an equally critical, yet underappreciated property, especially when switching costs are present. Systems that update policies too frequently may suffer from over-correction, chasing short-term forecast noise at the expense of long-term efficiency.
6 Conclusion
To summarize, the inclusion of switching costs in the FHC algorithm introduces a trade-off between forecast errors and switching costs. While traditional FHC formulations can penalize switching costs in the objective function, longer commitment periods provide a natural mechanism to use a stable forecast and reduce the need for frequent adjustments. This approach is more viable when optimization is computed over a set of scenarios, since the aggregation of policies offers a more robust policy, hedging against increasing forecast errors in a forecast horizon.
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. Our analysis shows that in systems where forecasts guide decision-making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs. Furthermore, we propose $SDC$, a novel metric to evaluate the stability of scenario sets.
Our study has several important limitations that should be acknowledged. First, our theoretical analysis assumes convex cost functions and exponentially decaying forecast errors, which may not fully capture the complexity of real-world energy systems. Second, while we demonstrate the relationship between forecast stability and optimization performance on a single case study, further validation across different energy management contexts and geographical locations would strengthen our findings. Additionally, our definition of switching costs primarily focuses on grid ramping, which serves as a proxy for the more complex switching costs in real energy systems, including wear-and-tear, efficiency losses, and network fees. Future work should address these limitations by testing across diverse datasets, and exploring more sophisticated methods for enhancing forecast stability without compromising accuracy. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape. It would also be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
Acknowledgments
This project has received funding from the European Union’s research and innovation programme Horizon Europe under the grant agreement No.101136211
Appendix A Cost Function and Optimization
A.1 Optimization score
Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization. The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. The optimization score is defined as:
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}}\right)$ (29)
When grid-related KPIs are included, the average includes the grid score $D$. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score. The average score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
The normalized electricity cost, $C$, as delineated in Eq. 31, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq. 32 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
$C=\frac{c_{{submission}}}{c_{{nobattery}}}$ (31)
$c=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times T_{h}\right)$ (32)
Similarly, the normalized carbon emissions, $G$, is defined in Eq. 33 as the ratio of district carbon emissions for a given policy, $g_{{policy}}$, relative to the emissions in the aforementioned baseline scenario, $g_{{nobattery}}$. The emission metric $g$ is elaborated in Eq. 34 as the sum of carbon emissions, measured in $\left({kg}_{\mathrm{CO}_{2}\mathrm{e}}/\mathrm{kWh}\right)$, given by $E_{h}\times O_{h}$. Here, $O_{h}$ represents the carbon intensity for the hour $h$.
$G=\frac{g_{{submission}}}{g_{{nobattery}}}$ (33)
$g=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times O_{h}\right)$ (34)
Lastly, the evaluation metric includes a grid-related KPI, $D$. The metric follows grid-level objectives, such as the minimization of ramping and load factor. It is defined as the mean of the normalized ramping KPI, $R$, and the Load Factor KPI, $(1-L)$. The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq. 35.
$D=\operatorname{avg}\left(\frac{R_{\text{ALG}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{ALG}}}{1-L_{\text{no\ battery}}}\right)$ (35)
The ramping KPI, $R$, reflects the smoothness of the district’s load profile. A low $R$ indicates a gradual increase in grid electricity demand even after self-generation becomes unavailable in the evening and early morning, while a high $R$ indicates abrupt changes in load on the grid, which may lead to unscheduled strain on grid infrastructure and potential blackouts due to supply deficits. It is calculated as the sum of the absolute difference of net electricity consumption between consecutive time steps:
$R=\sum_{t=1}^{8760}|E_{t}-E_{t-1}|$ (36)
The Load Factor, $L$, indicates the efficiency of electricity consumption and is bounded between 0 (very inefficient) and 1 (highly efficient). Thus, the goal is to minimize $(1-L)$. $L$ is calculated as the average ratio of monthly average to maximum net electricity consumption:
$L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (37)
A.2 Optimization Formulation
$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{S}_{tt^{\prime}})$ (38)
$=$ (39)
subject to
$SOC_{min}\leq SOC_{t}\leq SOC_{max},\quad\forall t,$ (40)
$x_{t}=x_{t}^{pos}+x_{t}^{neg},\quad\forall t,$ (41)
$-P_{max}\leq x_{t}^{neg}\leq 0,\quad\forall t,$ (42)
$SOC_{t}=SOC_{t-1}+$ (43)
$\eta_{charging}\cdot x_{t}^{pos}-\frac{1}{\eta_{discharging}}\cdot x_{t}^{neg},\quad\forall t.$ (44)
In this formulation:
  • $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the look-ahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process. We assume that the weight factors are equal for all scenarios. In the deterministic case, the number of scenarios is equal to 1.
  • $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively. Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
  • $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
  • $P_{max}$ is the maximum power with which the battery can be charged or discharged.
  • $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
A.3 In-sample Performance
Table 5: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours (with data from in-sample buildings)
Optimization Commitment Period ($v_{O}$)
Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
Panel A: Deterministic FHC with point forecast
1 0.486
2 0.489 0.489
3 0.491 0.492 0.491
4 0.492 0.492 0.494 0.492
5 0.493 0.494 0.494 0.494 0.493
6 0.493 0.493 0.493 0.495 0.495 0.493
7 0.493 0.494 0.495 0.495 0.495 0.495 0.493
8 0.493 0.493 0.495 0.493 0.495 0.496 0.495 0.493
9 0.494 0.495 0.494 0.495 0.495 0.495 0.495 0.496 0.494
10 0.494 0.494 0.495 0.495 0.494 0.495 0.496 0.495 0.496 0.494
11 0.494 0.495 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.494
12 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.495 0.496 0.496 0.495
Panel B: Stochastic FHC with probabilistic forecast
1 0.798
2 0.805 0.800
3 0.809 0.806 0.803
4 0.809 0.805 0.807 0.802
5 0.811 0.807 0.806 0.805 0.805
6 0.812 0.807 0.805 0.805 0.806 0.805
7 0.810 0.808 0.807 0.805 0.806 0.808 0.805
8 0.808 0.806 0.808 0.802 0.807 0.807 0.807 0.804
9 0.811 0.809 0.806 0.805 0.806 0.807 0.808 0.806 0.805
10 0.811 0.807 0.807 0.805 0.806 0.807 0.808 0.806 0.808 0.806
11 0.810 0.808 0.808 0.805 0.807 0.808 0.807 0.807 0.808 0.808 0.806
12 0.813 0.809 0.807 0.805 0.806 0.808 0.807 0.806 0.807 0.808 0.808 0.807
Note: Diagonal elements represent equal commitment periods for both forecast and optimization. Lower values indicate better performance. Empty cells (—) represent invalid combinations where $v_{O}>v_{F}$.
Accuracy
Stability: vertical (top) and horizontal (bottom)
Optimization Score: without switching costs
Optimization Score: with switching costs
Point Forecast
+ Deterministic FHC
Probabilistic forecast
+ Stochastic FHC
Figure 8: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from in-sample buildings.